OpenStudy (anonymous):

Prove the following \[\forall a,b \in \mathbb{r}, a < 0, b < 0 \implies a * b > 0\] Prove the following \[\forall a,b \in \mathbb{r}, a < 0, b < 0 \implies a * b > 0\] @Mathematics

jhonyy9 (jhonyy9):

because minus * minus is plus result that this is true allways

OpenStudy (anonymous):

I need a stronger proof

jhonyy9 (jhonyy9):

right ???

jhonyy9 (jhonyy9):

suppose that not is true and check the result so by reductio ad abbsurdum prove style, methode

OpenStudy (anonymous):

the book says you have to begin by first proving the following statement\[(-a) * b = -(a * b)\]

jhonyy9 (jhonyy9):

sorry

OpenStudy (anonymous):

(-a) * b + a*b = b(-a + a) = b*0 = 0, which we proved in the previous question hence, (-a)*b + a*b -(a*b) = -(a*b) therefore, (-a)*b = -(a*b)

OpenStudy (anonymous):

a<0, b<0 this implies -a> and -b>o this implies (-a)(-b)>0 this implies ab>0(because (-a)(-b)=ab)

OpenStudy (asnaseer):

ok, this is my attempt at this: you already proved that (-a)*b + ab = 0 therefore: -b*( (-a)*b + a*b ) = 0 (-b)*(-a)*b + (-b)*a*b = 0 ( (-b)*(-a) + (-b)*a )*b = 0 (-b)*(-a) + (-b)*a = 0 you also proved (-a)*b = -(a*b) therefore: (-b)*a = -(a*b) plug this into above to get: (-b)*(-a) + -(a*b) = 0 therefore: (-b)*(-a) = (a*b)

OpenStudy (anonymous):

@asnaseer that was how the book did it

OpenStudy (anonymous):

Consider a number \(c\) defined as \(c=ab+(-a)(b)+(-a)(-b)=ab+(-a)[b+(-b)]\). \(b+(-b)=0\). Therefor \(c=ab+(-a)[0] \implies c=ab\). We can also write \(c=b(a+(-a))+(-a)(-b) \implies c=b(0)+(-a)(-b)=(-a)(-b).\) We have \(c=ab\) and \(c=(-a)(-b)\), which implies that \(ab=(-a)(-b)\).